# Critical Nodes Identification in Complex Networks: A Survey

### v1

@@ 2.4 K-truss
where $\Delta_{ijk}$ represents a triangle with vertices $i$, $j$, and $k$, $k$ is a common neighbor of the two nodes, and $\Delta_{G}$ is the set of all triangles in network $G$. The higher the support of an edge, the higher its clustering characteristic. The decomposition proceeds based on edge support. Initially, set the parameter to $k=3$. Then, remove edges in the network with support less than $k-2$, and reduce the support of the two other edges forming a triangle with each edge by 1, updating the network topology. Continue removing edges with support less than $k-2$ until all edges in the network have support greater than or equal to $k-2$. Nodes removed in this stage are assigned to the k-truss layer. For the next step, increase $k+1$, and repeat the process of removing edges with support less than $k-2$ until all nodes in the network are completely assigned to the corresponding layers. Malliaros et al. malliaros2016locating introduced a triangle-graph-based k-core decomposition that refines the analysis of densely connected networks. Their method calculates the "trustness" of an edge, defined by its participation in a k-truss (the number of triangles it is part of). This approach offers a more detailed analysis compared to $k$-shell decomposition, making it particularly effective for identifying influential communities within networks.
@@ 2.6 Neighborhood Similarity
Degree Centrality RN2 RN3 RN4 ,  ma2019quasi zhu2024identifying $+$ Simple and computationally efficient. -- Fails to capture global network structure, only reflects local connectivity of neighbours.
$k$-shell Index RN41 RN42 ,  li2018identification ,  RN43 ,  zareie2018hierarchical ,  RN6 ,  liu2015core ,  liu2015improving $+$ Suitable for large-scale networks, identifies core nodes. -- Cannot differentiate nodes within the same layer, information spread may not correlate directly with $k$-shell values.
H-index lu2016h RN10 RN12 zareie2019ehc $+$ Incorporates degree and neighbor influence. -- Cannot distinguish nodes with the same H-index, dependent on neighborhood degree distribution.
K-truss 2008Trusses malliaros2016locating $+$ Identifies community cores and modules. -- High computational cost, ineffective in sparse networks.
Structural Hole yu2017critical yu2017identifying xu2018identifying liu2021identifying54 zhao2023ranking21 ma2023effective $+$ Identifies bridging nodes, effective for information spread. -- Computationally expensive, requires global information.
Neighborhood Similarity wang2015link lu2022critical11 ai2023identifying34 rao2022cbim tong2024novel $+$ Beneficial for social networks and recommendation systems. -- High computational complexity in large networks.
@@ 3.3 Pagerank Algorithm
Eigenvector Centrality bonacich1971factoring bonacich2007some ilyas2011identifying ,  PhysRevE71 ,  ahajjam2015leadersrank martin2014localization zhong2018identifying $+$ Considers neighbor importance, widely used in networks. -- Suffers from localization issues, slow convergence.
Katz Centrality katz1953new zhang2015identifying $+$ Balances local and global influence. -- High computational cost, sensitive to parameters.
PageRank Algorithm page1998pagerank yang2017mining hsu2017unsupervised ,  sheng2020identifying12 su2021identification67 $+$ Scales well for large networks, robust. -- Sensitive to damping factor, ignores node attributes.
Voting Mechanism RN33 ,  li2014identifying ,  RN36 RN38 ,  RN37 ,  li2022improved23 ,  liu2016evaluating $+$ Simple and adaptable for various networks. -- Affected by network structure, struggles in dense graphs.
@@ 4.2 Closeness Centrality
Betweenness Centrality RN56 ,  hage1995eccentricity ,  RN59 ,  RN61 ,  song2015novel ,  ventresca2015efficiently ,  zhang2011node ,  yang2020critical ,  kianian2021efficient ,  xiao2024new $+$ Captures shortest paths, global influence, robustness. -- High computational cost, sensitive to network changes.
Closeness Centrality freeman2002centrality latora2001efficient salavati2019ranking Okamoto7328 sheng2020identifying $+$ Measures reachability efficiency, effectiveness. -- Ineffective in disconnected networks, computationally expensive in large graphs.
Random Walk iannelli2018influencers RN7 kermarrec2011second $+$ Models information spread, considers network topology, improves influence ranking. -- Computationally intensive, rely on walk length.
@@ 5.2 Maximization of Connected Component
Network Control li2006controlling ,  ghosh2022synchronized ,  d2023controlling ,  liu2011controllability ,  ding2017key ,  lu2019pinning ,  Zhu10106394 ,  yu2012distributed ,  yu2013synchronization ,  ding2016optimizing ,  amani2017finding ,  amani2018controllability ,  liu2018optimizing ,  wang2012optimizing ,  bof2016role ,  zhou2018node ,  hui2021node ,  bomela2023finding ,  jiang2024pinning ,  sun2025optimal ,  zhang2017effect ,  mo2019effects ,  jiang2023searching ,  cao2024synchronization ,  zhang2022necessary ,  gao2025effects $+$ Beneficial for dynamic networks and control applications, quantifies node controllability. -- Requires accurate network topology information, relies on control theory.
Maximization of Connected Component Optimization aringhieri2016general ,  alozie2021efficient ,  Thulasiraman ,  Boginski0007 ,  Karygiannis1644271 ,  arulselvan2011cardinality ,  li2011finding ,  ventresca2015experimental ,  veremyev2014exact ,  ren2020identifying ,  wang2010community ,  lam2013identifying ,  ventresca2014fast ,  tang2020discrete $+$ Increases network fragmentation, applicable to security and IDS applications. -- Excessively breaks the network, affecting normal operations.
Minimization of Pair Connectivity Optimization di2012branch ,  ARULSELVAN20092193 ,  shen2012adaptive ,  ventresca2014derandomized ,  dinh2015assessing ,  sarker2019critical ,  purevsuren2016heuristic ,  addis2016hybrid ,  chen2020critical ,  ventresca2012global ,  shen2012discovery ,  ventresca2014region ,  yin2023mixed23 ,  zhang2024evolutionary ,  fortz2024min ,  jiang2025identifying ,  kouam2025novel ,  zhou2024finding $+$ Effective in suppressing spreading and network attacks. -- High computational demand in large-scale network.
Minimum Spanning Tree chen2004evaluation ,  di2011complexity ,  hermelin2016parameterized ,  addis2013identifying ,  aringhieri2016local ,  wang2017evaluation $+$ Simple and computationally efficient. -- cannot fully capture network complexity and dynamics.
@@ 5.3 Minimization of Pair Connectivity
Recent studies have focused on optimizing critical node detection and network robustness under adversarial conditions. Zhang et al. zhang2024evolutionary formulated the Critical Node Detection problem in interdependent networks as a Bi-objective Optimization Problem (BICND) and designed a method integrating single-layer and multi-layer network transfer. This approach reduced algorithmic complexity while enhancing accuracy in identifying critical nodes. Fortz et al. fortz2024min proposed a compact integer programming formulation for the node attack optimization problem using pseudo-components within a two-level model, addressing performance limitations of prior iterative methods. Jiang et al. jiang2025identifying introduced the EBC method, leveraging loop structures to assess network robustness against targeted attacks and precisely identify edges critical to resilience. Kouam et al. kouam2025novel simulated adversarial behavior to evaluate node influence, modeling attack progression toward the target. Zhou et al. zhou2024finding examined the NP-hard nature of the protection node problem and optimized key node selection via an influence redundancy mechanism. By minimizing the adjacency matrix’s spectral radius after key node removal, they demonstrated submodular properties in the objective function and achieved an approximation ratio of $O(N\log N)$. These advances enhance network resilience by refining critical node and edge, offering efficient solutions for targeted attacks. The multi-area power system can be described using power grid models and associated control models. In contrast, the networked $T-S$ fuzzy system is modeled using $T-S$ fuzzy models that consist of multiple fuzzy rules. The research on "Critical-metrics-based attack strategy and resilient $H_{\infty}$ state estimator design for multi-area power systems" presents a critical-metrics-based attack strategy. From the attacker’s perspective, this strategy targets the critical information packets in the key areas of the power system, thereby amplifying the impact of the attack on system performance. Correspondingly, from the defender’s perspective, a resilient state estimator is designed to alleviate the adverse effects of potential critical-metrics-based attacks.
@@ 6.3 Machine Learning
Graph Attention Network: Most existing node importance assessments in knowledge graphs fail to fully utilize available information and cannot capture node relationships and attribute features. To address this, Park et al. park2019estimating GENI uses predicate-aware attention mechanisms and flexible centrality adjustments to aggregate node and neighbor information across multiple GNN layers. It updates node embeddings via an aggregation function with a nonlinear transformation and maps these to importance scores using a scoring neural network. GENI also incorporates node in-degree centrality to adjust importance estimates. Define a GNN with $L$ layers of networks, where the $l$-th aggregates information from the $l-1$ th layer as follows,
@@
where $\overrightarrow{h}_{i}^{l-1}$ is the $v_{i}$ represents the information aggregation of the $(l-1)$ th layer, each neighbor node $N(i)$ of node $v_{i}$ denotes the set of its neighbors (each initialized with a weight $w_{i,j}^{l}$). The aggregation function and nonlinear transformation function share parameters across all nodes. At each layer, node scores are updated through a weighted combination of their own value and neighbors’ values, with attention mechanisms dynamically adjusting connection influence. Initial node scores come from feature embeddings processed through a neural scoring network. GENI also incorporates a centralization score based on node in-degree, refined through scaling and offset adjustments. Finally, a nonlinear activation function integrates this score with learned representations. Furthermore, Park et al. park2020multiimport designed MultiImport, an end-to-end latent variable model to infer node importance from multiple sparse and possibly overlapping input signals. MultiImport uses edge-aware attention aggregation and node centralization adjustment within a GNN framework to learn robust node representations. It models the knowledge graph structure and uses clustering to group nodes with similar functional signals. Munikoti et al. munikoti2022scalable introduced ILGR, a scalable GNN-based framework that approximates node and link importance through attention-driven embedding. ILGR encodes local structure by aggregating neighborhood information, with an attention mechanism assigning adaptive weights based on structural roles. Ge et al. ge2024node proposed MFA-NIE, integrating structural, relational, and attribute features with an enhanced attention mechanism for feature extraction, and adjusts node importance using TOPSIS centrality. Chen et al. chen2024social developed a hierarchical graph-enhanced variational autoencoder that reconstructs neighborhood representations to learn social influence patterns, using hierarchical graph attention networks to capture influential neighborhoods and mitigate data sparsity. Liu et al. liu2022nie23 introduced NIE-GAT for inter-domain routing networks, mapping static node features to dynamic cascading failure impacts through graph attention networks. Additionally, Kou et al. kou2023identify23 refined GCN-based approaches by framing node influence prediction as a regression task and using a multi-head attention mechanism to integrate neighbor importance, with the IC model providing simulation-based ground truth for optimization.
@@ 6.4 Reinforcement Learning
Information Entropy nikolaev2015efficient ,  zareie2019influential ,  nitt2016using ,  fu2015identifying ,  guo2020influential ,  xu2018discovery ,  tulu2018identifying ,  ai2017node ,  wu2023identify ,  tong2023vital ,  li2019key $+$ Robustness to various network structures. -- Dependent on accurate data, hard for large or dynamic networks.
Clustering Coefficient RN35 ,  gao2014ranking ,  yang2018identifying ,  zareie2020finding ,  dablander2019node ,  liu2022identifying11 ,  wang2022influential22 ,  zhang2013identifying $+$ Improves identification of influential nodes. -- Potential oversimplification of network dynamics, reliance on local information.
Graph Conventional Network ZHAO202018 ,  kumar2022influence23 ,  zhang2022new57 ,  liu2022learning $+$ Straightforward implementation, robustness in various network types. -- Limited ability to capture complex patterns, potential inefficiency with large-scale networks.
Graph Embeddings wei2024enimnr ,  keikha2020influence ,  bouyer2024maximizing ,  YihangandHu2024 ,  ahmad2024learning ,  rashid2024olapgn ,  xiong2024vital ,  yu2020identifying $+$ Improves accuracy for influential nodes. -- Reliance on quality of embeddings.
Graph Attention Network park2019estimating ,  park2020multiimport ,  munikoti2022scalable  ge2024node ,  chen2024social ,  liu2022nie23 ,  kou2023identify23 $+$ Heightened accuracy in ranking influential nodes, the ability to focus on important features. -- Dependency on hyperparameter tuning, potential overfitting in small datasets.
Graph Contrast Learning liu2023node ,  zhang2024label ,  shu2025node $+$ Effective identification of key nodes by leveraging differences between subgraph representations. -- High computational cost, potential overfitting.
Graph Neural Networks huang2022estimating ,  chen2024deep ,  lin2024node ,  zhao2023novel ,  michos2024critical $+$ Superior ability to capture complex patterns, adaptability to diverse network structures. -- Dependence on extensive training data and difficulties in model interpretability.
Reinforcement Learning fan2020finding ,  tan2024learning ,  jaques2019social ,  chen2023identifying79 ,  li2019disco ,  chen2023touplegdd ,  ling2023deep ,  li2022piano ,  uthayasuriyan2024adaptive ,  li2024adarisk ,  xu2025influence  zhu2025bigdn ,  ahmad2025learning $+$ Dynamic adaptation to network changes, improves decision-making over time. -- Complexity in setting reward structures, reliance on extensive exploration and training periods.

### v2

@@ 2.4 K-truss
where $\Delta_{ijk}$ represents a triangle with vertices $i$, $j$, and $k$, $k$ is a common neighbor of the two nodes, and $\Delta_{G}$ is the set of all triangles in network $G$. The higher the support of an edge, the higher its clustering characteristic. The decomposition proceeds based on edge support. Initially, set the parameter to $k=3$. Then, remove edges in the network with support less than k2k-2, and reduce the support of the two other edges forming a triangle with each edge by 1, updating the network topology. Continue removing edges with support less than k2k-2 until all edges in the network have support greater than or equal to k2k-2. Nodes removed in this stage are assigned to the k-truss layer. For the next step, increase $k+1$, and repeat the process of removing edges with support less than k2k-2 until all nodes in the network are completely assigned to the corresponding layers. Malliaros et al. malliaros2016locating introduced a triangle-graph-based k-core decomposition that refines the analysis of densely connected networks. Their method calculates the "trustness" of an edge, defined by its participation in a k-truss (the number of triangles it is part of). This approach offers a more detailed analysis compared to $k$-shell decomposition, making it particularly effective for identifying influential communities within networks.
@@ 2.6 Neighborhood Similarity
Degree Centrality RN2 RN3 RN4 ,  ma2019quasi zhu2024identifying ++ Simple and computationally efficient. - Fails to capture global network structure, only reflects local connectivity of neighbours.
$k$-shell Index RN41 RN42 ,  li2018identification ,  RN43 ,  zareie2018hierarchical ,  RN6 ,  liu2015core ,  liu2015improving ++ Suitable for large-scale networks, identifies core nodes. - Cannot differentiate nodes within the same layer, information spread may not correlate directly with $k$-shell values.
H-index lu2016h RN10 RN12 zareie2019ehc ++ Incorporates degree and neighbor influence. - Cannot distinguish nodes with the same H-index, dependent on neighborhood degree distribution.
K-truss 2008Trusses malliaros2016locating ++ Identifies community cores and modules. - High computational cost, ineffective in sparse networks.
Structural Hole yu2017critical yu2017identifying xu2018identifying liu2021identifying54 zhao2023ranking21 ma2023effective ++ Identifies bridging nodes, effective for information spread. - Computationally expensive, requires global information.
Neighborhood Similarity wang2015link lu2022critical11 ai2023identifying34 rao2022cbim tong2024novel ++ Beneficial for social networks and recommendation systems. - High computational complexity in large networks.
@@ 3.3 Pagerank Algorithm
Eigenvector Centrality bonacich1971factoring bonacich2007some ilyas2011identifying ,  PhysRevE71 ,  ahajjam2015leadersrank martin2014localization zhong2018identifying ++ Considers neighbor importance, widely used in networks. - Suffers from localization issues, slow convergence.
Katz Centrality katz1953new zhang2015identifying ++ Balances local and global influence. - High computational cost, sensitive to parameters.
PageRank Algorithm page1998pagerank yang2017mining hsu2017unsupervised ,  sheng2020identifying12 su2021identification67 ++ Scales well for large networks, robust. - Sensitive to damping factor, ignores node attributes.
Voting Mechanism RN33 ,  li2014identifying ,  RN36 RN38 ,  RN37 ,  li2022improved23 ,  liu2016evaluating ++ Simple and adaptable for various networks. - Affected by network structure, struggles in dense graphs.
@@ 4.2 Closeness Centrality
Betweenness Centrality RN56 ,  hage1995eccentricity ,  RN59 ,  RN61 ,  song2015novel ,  ventresca2015efficiently ,  zhang2011node ,  yang2020critical ,  kianian2021efficient ,  xiao2024new ++ Captures shortest paths, global influence, robustness. - High computational cost, sensitive to network changes.
Closeness Centrality freeman2002centrality latora2001efficient salavati2019ranking Okamoto7328 sheng2020identifying ++ Measures reachability efficiency, effectiveness. - Ineffective in disconnected networks, computationally expensive in large graphs.
Random Walk iannelli2018influencers RN7 kermarrec2011second ++ Models information spread, considers network topology, improves influence ranking. - Computationally intensive, rely on walk length.
@@ 5.2 Maximization of Connected Component
Network Control li2006controlling ,  ghosh2022synchronized ,  d2023controlling ,  liu2011controllability ,  ding2017key ,  lu2019pinning ,  Zhu10106394 ,  yu2012distributed ,  yu2013synchronization ,  ding2016optimizing ,  amani2017finding ,  amani2018controllability ,  liu2018optimizing ,  wang2012optimizing ,  bof2016role ,  zhou2018node ,  hui2021node ,  bomela2023finding ,  jiang2024pinning ,  sun2025optimal ,  zhang2017effect ,  mo2019effects ,  jiang2023searching ,  cao2024synchronization ,  zhang2022necessary ,  gao2025effects ++ Beneficial for dynamic networks and control applications, quantifies node controllability. - Requires accurate network topology information, relies on control theory.
Maximization of Connected Component Optimization aringhieri2016general ,  alozie2021efficient ,  Thulasiraman ,  Boginski0007 ,  Karygiannis1644271 ,  arulselvan2011cardinality ,  li2011finding ,  ventresca2015experimental ,  veremyev2014exact ,  ren2020identifying ,  wang2010community ,  lam2013identifying ,  ventresca2014fast ,  tang2020discrete ++ Increases network fragmentation, applicable to security and IDS applications. - Excessively breaks the network, affecting normal operations.
Minimization of Pair Connectivity Optimization di2012branch ,  ARULSELVAN20092193 ,  shen2012adaptive ,  ventresca2014derandomized ,  dinh2015assessing ,  sarker2019critical ,  purevsuren2016heuristic ,  addis2016hybrid ,  chen2020critical ,  ventresca2012global ,  shen2012discovery ,  ventresca2014region ,  yin2023mixed23 ,  zhang2024evolutionary ,  fortz2024min ,  jiang2025identifying ,  kouam2025novel ,  zhou2024finding ++ Effective in suppressing spreading and network attacks. - High computational demand in large-scale network.
Minimum Spanning Tree chen2004evaluation ,  di2011complexity ,  hermelin2016parameterized ,  addis2013identifying ,  aringhieri2016local ,  wang2017evaluation ++ Simple and computationally efficient. - cannot fully capture network complexity and dynamics.
@@ 5.3 Minimization of Pair Connectivity
Recent studies have focused on optimizing critical node detection and network robustness under adversarial conditions. Zhang et al. zhang2024evolutionary formulated the Critical Node Detection problem in interdependent networks as a Bi-objective Optimization Problem (BICND) and designed a method integrating single-layer and multi-layer network transfer. This approach reduced algorithmic complexity while enhancing accuracy in identifying critical nodes. Fortz et al. fortz2024min proposed a compact integer programming formulation for the node attack optimization problem using pseudo-components within a two-level model, addressing performance limitations of prior iterative methods. Jiang et al. jiang2025identifying introduced the EBC method, leveraging loop structures to assess network robustness against targeted attacks and precisely identify edges critical to resilience. Kouam et al. kouam2025novel simulated adversarial behavior to evaluate node influence, modeling attack progression toward the target. Zhou et al. zhou2024finding examined the NP-hard nature of the protection node problem and optimized key node selection via an influence redundancy mechanism. By minimizing the adjacency matrix’s spectral radius after key node removal, they demonstrated submodular properties in the objective function and achieved an approximation ratio of $O(N\log N)$. These advances enhance network resilience by refining critical node and edge, offering efficient solutions for targeted attacks. The multi-area power system can be described using power grid models and associated control models. In contrast, the networked TST-S fuzzy system is modeled using TST-S fuzzy models that consist of multiple fuzzy rules. The research on "Critical-metrics-based attack strategy and resilient $H_{\infty}$ state estimator design for multi-area power systems" presents a critical-metrics-based attack strategy. From the attacker’s perspective, this strategy targets the critical information packets in the key areas of the power system, thereby amplifying the impact of the attack on system performance. Correspondingly, from the defender’s perspective, a resilient state estimator is designed to alleviate the adverse effects of potential critical-metrics-based attacks.
@@ 6.3 Machine Learning
Graph Attention Network: Most existing node importance assessments in knowledge graphs fail to fully utilize available information and cannot capture node relationships and attribute features. To address this, Park et al. park2019estimating GENI uses predicate-aware attention mechanisms and flexible centrality adjustments to aggregate node and neighbor information across multiple GNN layers. It updates node embeddings via an aggregation function with a nonlinear transformation and maps these to importance scores using a scoring neural network. GENI also incorporates node in-degree centrality to adjust importance estimates. Define a GNN with $L$ layers of networks, where the $l$-th aggregates information from the l1l-1 th layer as follows,
@@
where $\overrightarrow{h}_{i}^{l-1}$ is the $v_{i}$ represents the information aggregation of the (l1)(l-1) th layer, each neighbor node $N(i)$ of node $v_{i}$ denotes the set of its neighbors (each initialized with a weight $w_{i,j}^{l}$). The aggregation function and nonlinear transformation function share parameters across all nodes. At each layer, node scores are updated through a weighted combination of their own value and neighbors’ values, with attention mechanisms dynamically adjusting connection influence. Initial node scores come from feature embeddings processed through a neural scoring network. GENI also incorporates a centralization score based on node in-degree, refined through scaling and offset adjustments. Finally, a nonlinear activation function integrates this score with learned representations. Furthermore, Park et al. park2020multiimport designed MultiImport, an end-to-end latent variable model to infer node importance from multiple sparse and possibly overlapping input signals. MultiImport uses edge-aware attention aggregation and node centralization adjustment within a GNN framework to learn robust node representations. It models the knowledge graph structure and uses clustering to group nodes with similar functional signals. Munikoti et al. munikoti2022scalable introduced ILGR, a scalable GNN-based framework that approximates node and link importance through attention-driven embedding. ILGR encodes local structure by aggregating neighborhood information, with an attention mechanism assigning adaptive weights based on structural roles. Ge et al. ge2024node proposed MFA-NIE, integrating structural, relational, and attribute features with an enhanced attention mechanism for feature extraction, and adjusts node importance using TOPSIS centrality. Chen et al. chen2024social developed a hierarchical graph-enhanced variational autoencoder that reconstructs neighborhood representations to learn social influence patterns, using hierarchical graph attention networks to capture influential neighborhoods and mitigate data sparsity. Liu et al. liu2022nie23 introduced NIE-GAT for inter-domain routing networks, mapping static node features to dynamic cascading failure impacts through graph attention networks. Additionally, Kou et al. kou2023identify23 refined GCN-based approaches by framing node influence prediction as a regression task and using a multi-head attention mechanism to integrate neighbor importance, with the IC model providing simulation-based ground truth for optimization.
@@ 6.4 Reinforcement Learning
Information Entropy nikolaev2015efficient ,  zareie2019influential ,  nitt2016using ,  fu2015identifying ,  guo2020influential ,  xu2018discovery ,  tulu2018identifying ,  ai2017node ,  wu2023identify ,  tong2023vital ,  li2019key ++ Robustness to various network structures. - Dependent on accurate data, hard for large or dynamic networks.
Clustering Coefficient RN35 ,  gao2014ranking ,  yang2018identifying ,  zareie2020finding ,  dablander2019node ,  liu2022identifying11 ,  wang2022influential22 ,  zhang2013identifying ++ Improves identification of influential nodes. - Potential oversimplification of network dynamics, reliance on local information.
Graph Conventional Network ZHAO202018 ,  kumar2022influence23 ,  zhang2022new57 ,  liu2022learning ++ Straightforward implementation, robustness in various network types. - Limited ability to capture complex patterns, potential inefficiency with large-scale networks.
Graph Embeddings wei2024enimnr ,  keikha2020influence ,  bouyer2024maximizing ,  YihangandHu2024 ,  ahmad2024learning ,  rashid2024olapgn ,  xiong2024vital ,  yu2020identifying ++ Improves accuracy for influential nodes. - Reliance on quality of embeddings.
Graph Attention Network park2019estimating ,  park2020multiimport ,  munikoti2022scalable  ge2024node ,  chen2024social ,  liu2022nie23 ,  kou2023identify23 ++ Heightened accuracy in ranking influential nodes, the ability to focus on important features. - Dependency on hyperparameter tuning, potential overfitting in small datasets.
Graph Contrast Learning liu2023node ,  zhang2024label ,  shu2025node ++ Effective identification of key nodes by leveraging differences between subgraph representations. - High computational cost, potential overfitting.
Graph Neural Networks huang2022estimating ,  chen2024deep ,  lin2024node ,  zhao2023novel ,  michos2024critical ++ Superior ability to capture complex patterns, adaptability to diverse network structures. - Dependence on extensive training data and difficulties in model interpretability.
Reinforcement Learning fan2020finding ,  tan2024learning ,  jaques2019social ,  chen2023identifying79 ,  li2019disco ,  chen2023touplegdd ,  ling2023deep ,  li2022piano ,  uthayasuriyan2024adaptive ,  li2024adarisk ,  xu2025influence  zhu2025bigdn ,  ahmad2025learning ++ Dynamic adaptation to network changes, improves decision-making over time. - Complexity in setting reward structures, reliance on extensive exploration and training periods.